OpenStudy (anonymous):

Carbon 14 has a half-life of 1370 years. We start with 100% of Carbon 14. After how many years will there be 30% of Carbon 14 left?

OpenStudy (anonymous):

\[.3=(\frac{1}{2})^{\frac{t}{1370}}\] solve for t

OpenStudy (anonymous):

step one \[\frac{\ln(.3)}{\ln(.5)}=\frac{t}{1370}\] step two \[t=\frac{1375\times \ln(.3)}{\ln(.5)}\] step 4 use a calculator

OpenStudy (anonymous):

i'm just wondering, why do you use ln as opposed to log?

OpenStudy (anonymous):

makes no difference

OpenStudy (anonymous):

all logs are the same, change of base formula tells you that

OpenStudy (anonymous):

all right, thank you.

OpenStudy (anonymous):

yw